OpenStudy (anonymous):

Hey! If f(x) = -1/2x -3 find f^1(x)

sam (.sam.):

f '(x)?

OpenStudy (anonymous):

Yeah f(x)

sam (.sam.):

2/(2x+3)^2

OpenStudy (ash2326):

We have \[ f(x)= -\frac{1}{2x-3}\] Let's substitute \[x=f^{-1} (x)\] We get \[ x= -\frac{1}{2f^{-1} (x)-3}\] Let's cross multiply \[2xf^{-1} (x)-3x=-1\] Add 3x to both sides \[2xf^{-1} (x)=3x-1\] we get \[ f^{-1} (x)= \frac{3x-1}{2x}\]

sam (.sam.):

I think she's asking for derivative.

OpenStudy (mertsj):

What do you need to find, Lib, the first derivative or the inverse?

OpenStudy (anonymous):

Yessir

OpenStudy (anonymous):

I would appreciate the help